# Best Ergodic Averages via Optimal Graph Filters in Reversible Markov Chains

### v1

@@ III Graph Signal Processing
$TV_{f}\coloneqq\frac{\left(\sum_{x\in{\mathsf{X}}}\sum_{x\rightarrow y}\pi(x)P(x,y)|f(x)-f(y)|^{2}\right)^{1/2}}{\left(\sum_{x\in{\mathsf{X}}}\pi(x)|f(x)|^{2}\right)^{1/2}}.$
In this definition, the denominator serves as a normalization constant to obtain a dimensionless quantity. The numerator quantifies the weighted difference of signals on vertices from their neighbors, with weights provided by the matrix $DP$. The value $\pi(x)P(x,y)$ represents how similar the vertices $x$ and $y$ are. Here, $\pi(x)$ indicates the importance of vertex $x$ in terms of the stationary distribution. If $\pi(x)\approx 0$, vertex $x$ holds little importance, as the number of visits to this vertex in the Markov chain is very low compared to other vertices according to the ergodic theorem. Additionally, $P(x,y)$ quantifies the (directed) similarity between vertex $x$ and $y$ or it can be interpreted as the importance of $y$ to $x$ because it is proportional to the number of transitions from vertex $x$ to $y$. Thus, the term $\pi(x)P(x,y)|f(x)-f(y)|^{2}$ captures the importance of vertex $x$ multiplied by the weighted variation of the graph signal $f$ from $x$ to $y$, with the significance of this variation determined by $P(x,y)$.
@@
Proof.
The proof of this lemma is very similar to [15, Remark 3.1].
For any $x\in{\mathsf{X}}$, we can write $Lf(x)=f(x)-\sum_{x\rightarrow y}P(x,y)f(y)=\sum_{x\rightarrow y}P(x,y)(f(x)-f(y))$ as $\sum_{x\rightarrow y}P(x,y)=1$. Hence,
$2\langle f,Lf\rangle_{\pi}$
$=\sum_{x\in{\mathsf{X}}}\sum_{x\rightarrow y}\pi(x)P(x,y)f(x)(f(x)-f(y))$
$\phantom{xxxxx}+\sum_{y\in{\mathsf{X}}}\sum_{y\rightarrow x}\pi(y)P(y,x)f(y)(f(y)-f(x))$
$=\sum_{x\in{\mathsf{X}}}\bigg{(}\sum_{x\rightarrow y}\pi(x)P(x,y)f(x)(f(x)-f(y))$
$\phantom{xxxxx}+\sum_{y\rightarrow x}\pi(y)P(y,x)f(y)(f(y)-f(x))\bigg{)}$
as $x\rightarrow y$ if and only if $y\rightarrow x$ by reversibility. Since $\pi(x)P(x,y)=\pi(y)P(y,x)$, we have
$\pi(y)P(y,x)f(y)=\pi(x)P(x,y)f(y).$
Hence,
$\sum_{x\rightarrow y}\pi(x)P(x,y)f(x)(f(x)-f(y))$
$\phantom{xxxxx}+\sum_{y\rightarrow x}\pi(y)P(y,x)f(y)(f(y)-f(x))$
@@
This implies that
$2\langle f,Lf\rangle_{\pi}=\sum_{x\in{\mathsf{X}}}\sum_{x\rightarrow y}\pi(x)P(x,y)(f(x)-f(y))^{2}=TV_{f}^{2}$
as $\|f\|_{\pi}=1$. This completes the proof. ∎
This result implies that $\langle f,Lf\rangle_{\pi}$ is high when the variation of $f$ is high, and low when the variation of $f$ is low. Therefore, we can use $\langle f,Lf\rangle_{\pi}$ to quantify the variation of graph signal $f$.
Let us write the eigenvalues of $L$ as $\{\lambda_{x}\}_{x\in{\mathsf{X}}}\subset[0,2]$, where one of them is zero, say $\lambda_{x_{*}}=0$. These eigenvalues have corresponding real eigenfunctions $\{f_{x}\}_{x\in{\mathsf{X}}}$, which form an orthonormal basis for the space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ and $f_{x_{*}}={\bf 1}$. For any eigenfunction $f_{x}$, the following holds:
@@
Definition 4.
For any vertex $x\in{\mathsf{X}}$, its $k$-hop neighborhood, ${\mathcal{N}}_{k}(x)$, is the set all vertices that are part of a directed path with no more than $k$ edges starting from $x$. That is, $y\in{\mathcal{N}}_{k}(x)$ if there exists a directed path $\{x_{0},x_{1},\ldots,x_{l}\}$, where $l\leq k$, $x_{0}=x$, and $x_{l}=y$, such that $P(x_{i},x_{i+1})>0$ for all $i=0,\ldots,l-1$.
Now, it is time to give the definition of a graph filter.
Definition 5.
@@
In graph signal processing, the linear graph filter $H$ is considered to be local at $x$ if $Hf(x)$ only depends on $k$-hop neighbors of $x$, where $|{\mathcal{N}}_{k}(x)|\ll|{\mathsf{X}}|$. For instance,
$Pf(x)=\sum_{y\in{\mathsf{X}}}f(y)P(x,y)=\sum_{y\in{\mathcal{N}}_{1}(x)}f(y)P(x,y);$
that is, to compute $Pf(x)$, we need $1$-hop neighbors of $x$. Therefore, $P$ is a local graph filter. Similarly, $L$ is also a local graph filter. Indeed, $P$ and $L$ are called one-hop graph filters due to the following definition.
Definition 6.
A graph filter $H$ is called $k$-hop filter if $Hf(x)$ can be computed via the elements of ${\mathcal{N}}_{k}(x)$ for all $x\in{\mathsf{X}}$.
@@
Let $Z$ represent a generic one-hop filter, such as $Z=P$, $Z=L$, or $Z=A$, among others. In graph signal processing, $Z$ serves as a generalization of the shift operator used in classical signal processing. For any positive integer $k$, $Z^{k}$ is a $k$-hop filter. Initially, our definition of a graph filter required only that the operator $H$ be linear, allowing any arbitrary linear operator to be chosen for $H$. However, we now focus on filters that can be expressed as polynomials of $Z$.
Definition 7.
@@
Given a one-hop operator $Z$, the polynomial graph filter $H$ is a linear operator of the following form $H=p(Z)$ for some polynomial $p$ of some degree $K$. Hence,
$H=\sum_{k=0}^{K}a_{k}Z^{k},$
where $Z^{0}=\mathbb{Id}$.
@@
To design a polynomial filter, one should first select a specific one-hop filter $Z$. In our case, this one-hop filter will be the combinatorial Laplacian $L$ since we have a precise interpretation of its eigenvalues as frequencies. Recall that the eigenvalues of $L$ are $\{\lambda_{x}\}_{x\in{\mathsf{X}}}\subset[0,2]$, where $\lambda_{x_{*}}=0$, with the corresponding real eigenfunctions $\{f_{x}\}_{x\in{\mathsf{X}}}$ forming an orthonormal basis for the space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle{\pi})$ and $f_{x_{*}}=\mathbf{1}$. Thus, any graph signal $f$ has the following unique representation:
$f=\sum_{x\in{\mathsf{X}}}\langle f,f_{x}\rangle_{\pi}f_{x}.$
This representation leads to the definition of graph Fourier transform.
Definition 8.
[…]
@@
Obviously $\mathcal{L}$ is a linear operator from $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ to $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle)$. Moreover, for any $x\in{\mathsf{X}}$, we have $\mathcal{L}(f_{x})=e_{x}$, where $e_{x}(y)\coloneqq 1_{\{x=y\}}$. Hence, $\mathcal{L}$ maps orthonormal basis $\{f_{x}\}_{x\in{\mathsf{X}}}$ of $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ to orhonormal basis $\{e_{x}\}_{x\in{\mathsf{X}}}$ of $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle)$. Therefore, it is a unitary transformation. Hence, the inverse of the GFT $\mathcal{L}$ is its adjoint: $\mathcal{L}^{-1}=\mathcal{L}^{*}$, where $\mathcal{L}^{*}$ is the unique operator from $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle)$ to $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ satisfying the following:
@@
Let $H$ be a polynomial graph filter of degree $K$; that is, $H=\sum_{k=0}^{K}a_{k}L^{k}\eqqcolon p(L)$. This filter then acts on any graph signal $f$ with GFT ${\hat{f}}$ in the following manner:
@@
${\hat{H}}\coloneqq\mathop{\rm diag}(\{p(\lambda_{x})\}_{x\in{\mathsf{X}}})=p(\Lambda),$
where $\Lambda\coloneqq\mathop{\rm diag}(\{\lambda_{x}\}_{x\in{\mathsf{X}}})$; that is, $\Lambda$ is the diagonal matrix containing the eigenvalues of $L$ as its diagonal entries. Given ${\hat{H}}$, we can describe the action of the polynomial filter $H$ in the frequency domain as follows:
${\hat{H}}({\hat{f}})=\mathop{\rm diag}(\{{\hat{f}}(x)\,p(\lambda_{x})\}_{x\in{\mathsf{X}}})=\mathcal{L}(H(f)).$
Unfortunately, since $\pi$ is generally unavailable, it is in general unfeasible to compute the GFT of any graph signal $f$ and frequency representation of any polynomial graph filter $H$. Consequently, filtering operations must be performed in the node domain instead of the frequency domain. Nevertheless, we can still gain insight into designing graph filters in the node domain by examining their effects in the frequency domain.
@@
Informal Definition of the Problem: For each degree $K$, find the optimal polynomial filter $H_{K}$ with the corresponding polynomial $p_{K}$ such that $H_{K}(f)$ converges to $\pi(f){\bf 1}$ as quickly as possible. To achieve this convergence, the polynomial $p_{K}$ should attenuate the eigenvalues (or frequencies) other than 00 of $L$. Therefore, the polynomial graph filter should function as a low-pass filter.
Revisiting Ergodic Theorem
Note that the iteration in the ergodic theorem can be written as a polynomial of $P$ in the following form:
$\frac{(\mathbb{Id}+P+\ldots+P^{t-1})}{t}\eqqcolon p_{t}(P),$
where
$p_{t}(z)\coloneqq\sum_{k=0}^{t-1}\frac{1}{t}z^{t}=\frac{1}{t}\frac{1-z^{t}}{1-z},\,\,\,z\in[-1,1).$
If we define the following polynomial $q_{t}(z)\coloneqq p_{t}(1-z)$, then we can write the same expression in terms of $L$ as follows:
$p_{t}(P)$ $=q_{t}(L)=\frac{1}{t}\frac{1-(1-L)^{t}}{1-(1-L)}$
$=\frac{1}{t}\frac{1-\sum_{k=0}^{t}{t\choose k}(-L)^{k}}{L}$
$=\sum_{k=1}^{t}\frac{1}{t}{t\choose k}(-1)^{k-1}L^{k-1}$
$\eqqcolon\sum_{k=0}^{t-1}a_{e}(k)\,L^{k}.$
If we define the polynomial graph filter $H_{t}\coloneqq q_{t}(L)$, then we observe that we are indeed applying the polynomial graph filter $H_{t}$ to the graph signal $f$ in each iteration of the ergodic theorem. As $t$ approaches infinity, this graph filter attenuates the frequencies different from zero. Specifically,
@@
for $z\in(0,2]$. This can be proved very easily since $p_{t}(z)\rightarrow 0$ as $t\rightarrow\infty$ for $z\in[-1,1)$. Consequently, $H_{t}(f)\rightarrow\pi(f){\bf 1}$. Hence, in each iteration of the ergodic theorem, we are applying a low-pass filter to the graph signal $f$. However, this low-pass filter is not the optimal one. The goal of this paper is to find or approximate the optimal low-pass filter in each iteration. $\square$
Given the non-zero eigenvalues $\{\lambda\}_{x\in{\mathsf{X}}\setminus\{x_{*}\}}\subset(0,2]$ of $L$, it is not always the case that $q_{t}(\lambda_{x})=0$ for all $x\in{\mathsf{X}}\setminus\{x_{*}\}$ for sufficiently large $t$ values, where $q_{t}$ is the polynomial used in the ergodic theorem. Indeed, if we have complete knowledge of the eigenvalues $\{\lambda\}_{x\in{\mathsf{X}}}$ of $L$, then we can use a Lagrange polynomial $q$ of degree $|{\mathsf{X}}|-1$ [16, Definition 6.1] to achieve:
@@
However, in general, we do not have access to the full set of eigenvalues of $L$ and so constructing $q_{t}^{L}$ is unfeasible. In general, the most that one can do is to estimate some important eigenvalues different than $\lambda_{x_{*}}$. For instance, we may obtain some lower bound to the second smallest eigenvalue of $L$ [17] [18, Section 9.2]. Indeed, if $\lambda_{x_{**}}$ denotes the second smallest eigenvalue of $L$, then $\lambda_{x_{**}}$ is equal to the spectral gap of $P$. In this case, if we have some lower bound on this second smallest eigenvalue:
$\lambda_{x_{**}}\geq\lambda_{{\mathsf{low}}},$
then the best approach would be to find a polynomial that satisfies:
@@
Here, we want to converge to zero for values in $[\lambda_{{\mathsf{low}}},2]$ as quickly as possible through polynomials of the form $q_{t}(z)=\sum_{k=0}^{t-1}a(k)z^{k}$ while maintaining $q_{t}(\lambda_{x_{*}})=1$. Thus, we can interpret our problem as designing an optimal (in some sense) polynomial low-pass graph filter with one-hop operator $L$ to filter out frequencies greater than $\lambda_{{\mathsf{low}}}$. In the following, we assume that $\lambda_{{\mathsf{low}}}$ is fixed. Using this assumption, we provide a precise mathematical description of the optimization problem we aim to solve.
@@
$\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\coloneqq\sup_{z\in[\lambda_{{\mathsf{low}}},2]}|p(z)|.$
Let us now give a motivation for this optimization problem. Note that, originally, we want to minimize $\|p_{t}(L)f-\pi(f){\bf 1}\|_{\pi}$ for any unit norm graph signal $f$, where $p_{t}\in P_{t-1}[\lambda_{{\mathsf{low}}},2]$ with $p_{t}(0)=1$, for each time step $t\geq 2$ as any graph signal is viewed as an element of the inner product space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ [^1] In this context, we can use any norm we prefer because the interpretation at the end will be the same.. An upper bound to this term can be obtained as follows:
$\|p_{t}(L)f-\pi(f){\bf 1}\|_{\pi}$ $=\left\|\sum_{x\in{\mathsf{X}}}{\hat{f}}(x)p_{t}(\lambda_{x})f_{x}-{\hat{f}}(x_{*})f_{x_{*}}\right\|_{\pi}$
$=\left\|\sum_{x\neq x_{*}}{\hat{f}}(x)p_{t}(\lambda_{x})f_{x}\right\|_{\pi}$
$\leq\sum_{x\neq x_{*}}|{\hat{f}}(x)||p_{t}(\lambda_{x})|\left\|f_{x}\right\|_{\pi}$
$=\sum_{x\neq x_{*}}|{\hat{f}}(x)||p_{t}(\lambda_{x})|$
$\leq\left(\sum_{x\neq x_{*}}|{\hat{f}}(x)|\right)\sup_{x\neq x_{*}}|p_{t}(\lambda_{x})|$
In the last inequality above, the first term is fixed. Thus, our objective should be to minimize the second term by choosing $p_{t}$ appropriately. However, because we do not have access to the full eigenvalue structure, we cannot directly minimize the second term. Given the lower bound $\lambda_{{\mathsf{low}}}$ for the second smallest eigenvalue of $L$, we can instead try to minimize $\|p_{t}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}$, which leads to the optimization problem ${\bf(OPT_{t})}$.
IV Filter Design with Bernstein Polynomials
For each time step $t\geq 2$, we want to design a polynomial $p$ of degree at most $t-1$ so that

### v2

@@ 3 Graph Signal Processing Approach to Birkhoff Ergodic Theorem
$TV_{f}\triangleq\frac{\left(\sum_{x\in{\mathsf{X}}}\sum_{x\rightarrow y}\pi(x)P(x,y)|f(x)-f(y)|^{2}\right)^{1/2}}{\left(\sum_{x\in{\mathsf{X}}}\pi(x)|f(x)|^{2}\right)^{1/2}}.$
In this definition, the denominator serves as a normalization constant to obtain a dimensionless quantity. The numerator quantifies the weighted difference of signals on vertices from their neighbors, with weights provided by the matrix $DP$, where $D\triangleq\mathop{\rm diag}(\pi)$. The value $\pi(x)P(x,y)$ represents how similar the vertices $x$ and $y$ are. Here, $\pi(x)$ indicates the importance of vertex $x$ in terms of the stationary distribution. If $\pi(x)\approx 0$, vertex $x$ holds little importance, as the number of visits to this vertex in the Markov chain is very low compared to other vertices according to the ergodic theorem. Additionally, $P(x,y)$ quantifies the (directed) similarity between vertex $x$ and $y$ or it can be interpreted as the importance of $y$ to $x$ because it is proportional to the number of transitions from vertex $x$ to $y$. Thus, the term $\pi(x)P(x,y)|f(x)-f(y)|^{2}$ captures the importance of vertex $x$ multiplied by the weighted variation of the graph signal $f$ from $x$ to $y$, with the significance of this variation determined by directed similarity $P(x,y)$.
@@
Proof 3.4.
The proof of this lemma is very similar to [14, Remark 3.1]. For any $x\in{\mathsf{X}}$, we can write $Lf(x)=f(x)-\sum_{x\rightarrow y}P(x,y)f(y)=\sum_{x\rightarrow y}P(x,y)(f(x)-f(y))$ as $\sum_{x\rightarrow y}P(x,y)=1$. Hence,
$2\langle f,Lf\rangle_{\pi}$ $=\sum_{x\in{\mathsf{X}}}\sum_{x\rightarrow y}\pi(x)P(x,y)f(x)(f(x)-f(y))+\sum_{y\in{\mathsf{X}}}\sum_{y\rightarrow x}\pi(y)P(y,x)f(y)(f(y)-f(x))$
$=\sum_{x\in{\mathsf{X}}}\bigg(\sum_{x\rightarrow y}\pi(x)P(x,y)f(x)(f(x)-f(y))+\sum_{y\rightarrow x}\pi(y)P(y,x)f(y)(f(y)-f(x))\bigg)$
as $x\rightarrow y$ if and only if $y\rightarrow x$ by reversibility. Since $\pi(x)P(x,y)=\pi(y)P(y,x)$, we have $\pi(y)P(y,x)f(y)=\pi(x)P(x,y)f(y).$ Hence,
$\sum_{x\rightarrow y}\pi(x)P(x,y)f(x)(f(x)-f(y))$ $+\sum_{y\rightarrow x}\pi(y)P(y,x)f(y)(f(y)-f(x))$
@@
This implies that $2\langle f,Lf\rangle_{\pi}=\sum_{x\in{\mathsf{X}}}\sum_{x\rightarrow y}\pi(x)P(x,y)(f(x)-f(y))^{2}=TV_{f}^{2}$ as $\|f\|_{\pi}=1$.
This result implies that $\langle f,Lf\rangle_{\pi}$ is high when the variation of $f$ is high, and low when the variation of $f$ is low. Therefore, we can use $\langle f,Lf\rangle_{\pi}$ to quantify the variation of graph signal $f$. Let us write the eigenvalues of $L$ as $\{\lambda_{x}\}_{x\in{\mathsf{X}}}\subset[0,2]$, where one of them is zero, say $\lambda_{x_{*}}=0$. These eigenvalues have corresponding real eigenfunctions $\{f_{x}\}_{x\in{\mathsf{X}}}$, which form an orthonormal basis for the space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ and $f_{x_{*}}={\bf 1}$. For any eigenfunction $f_{x}$, the following holds:
@@
3.1 Graph Filters
For any vertex $x\in{\mathsf{X}}$, its $k$-hop neighborhood, ${\mathcal{N}}_{k}(x)$, is the set all vertices that are part of a directed path with no more than $k$ edges starting from $x$. Now, it is time to give the definition of a graph filter.
Definition 3.5.
@@ 3.1 Graph Filters
An operator $H$ on graph signals is called $k$-hop operator if $Hf(x)$ can be computed via the elements of ${\mathcal{N}}_{k}(x)$ for all $x\in{\mathsf{X}}$. Hence, $P$ and $L$ are one-hop operators. Let $Z$ represent a generic one-hop operator, such as $Z=P$, $Z=L$, or $Z=A$, among others. In graph signal processing, $Z$ serves as a generalization of the shift operator used in classical signal processing. For any positive integer $k$, $Z^{k}$ is a $k$-hop operator. Initially, our definition of a graph filter required only that the operator $H$ be linear, allowing any arbitrary linear operator to be chosen for $H$. However, we now focus on filters that can be expressed as polynomials of $Z$.
Definition 3.6.
@@
Given a one-hop operator $Z$, the polynomial graph filter $H$ is a linear operator of the following form $H=p(Z)$ for some polynomial $p$ of some degree $K$. Hence, $H=\sum_{k=0}^{K}a_{k}Z^{k},$ where $Z^{0}=\mathbb{Id}$.
@@
To design a polynomial filter, one should first select a specific one-hop operator $Z$. In our case, this one-hop operator will be the combinatorial Laplacian $L$ since we have a precise interpretation of its eigenvalues as frequencies. Recall that the eigenvalues of $L$ are $\{\lambda_{x}\}_{x\in{\mathsf{X}}}\subset[0,2]$, where $\lambda_{x_{*}}=0$, with the corresponding real eigenfunctions $\{f_{x}\}_{x\in{\mathsf{X}}}$ forming an orthonormal basis for the space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle{\pi})$ and $f_{x_{*}}=\mathbf{1}$. Thus, any graph signal $f$ has the following unique representation: $f=\sum_{x\in{\mathsf{X}}}\langle f,f_{x}\rangle_{\pi}f_{x}.$ This representation leads to the definition of graph Fourier transform.
Definition 3.7.
The vector ${\hat{f}}\triangleq(\langle f,f_{x}\rangle_{\pi})_{x\in{\mathsf{X}}}$ in $\mathbb{R}^{{\mathsf{X}}}$, where $\mathbb{R}^{{\mathsf{X}}}$ is now endowed with the usual Euclidean inner product $\langle\cdot,\cdot\rangle$, is defined as the graph Fourier transform (GFT) of the graph signal $f$. Let us denote the GFT operator via $\mathcal{L}$ that performs this operation, and so, $\mathcal{L}(f)={\hat{f}}$.
@@
Obviously $\mathcal{L}$ is a linear operator from $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ to $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle)$. Moreover, for any $x\in{\mathsf{X}}$, we have $\mathcal{L}(f_{x})=e_{x}$, where $e_{x}(y)\triangleq 1_{\{x=y\}}$. Hence, $\mathcal{L}$ maps orthonormal basis $\{f_{x}\}_{x\in{\mathsf{X}}}$ of $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ to orhonormal basis $\{e_{x}\}_{x\in{\mathsf{X}}}$ of $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle)$. Therefore, it is a unitary transformation. Hence, the inverse of the GFT $\mathcal{L}$ is its adjoint: $\mathcal{L}^{-1}=\mathcal{L}^{*}$, where $\mathcal{L}^{*}$ is the unique operator from $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle)$ to $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ satisfying the following:
@@
Let $H$ be a polynomial graph filter of degree $K$; that is, $H=\sum_{k=0}^{K}a_{k}L^{k}\triangleq p(L)$. This filter then acts on any graph signal $f$ with GFT $^{\hat{f}}$ in the following manner:
@@
${\hat{H}}\triangleq\mathop{\rm diag}(\{p(\lambda_{x})\}_{x\in{\mathsf{X}}})=p(\Lambda),$
where $\Lambda\triangleq\mathop{\rm diag}(\{\lambda_{x}\}_{x\in{\mathsf{X}}})$; that is, $\Lambda$ is the diagonal matrix containing the eigenvalues of $L$ as its diagonal entries. Given $^{\hat{H}}$, we can describe the action of the polynomial filter $H$ in the frequency domain as follows: ${\hat{H}}({\hat{f}})=\mathop{\rm diag}(\{{\hat{f}}(x)\,p(\lambda_{x})\}_{x\in{\mathsf{X}}})=\mathcal{L}(H(f)).$ Unfortunately, since $\pi$ is generally unavailable, it is in general unfeasible to compute the GFT of any graph signal $f$ and frequency representation of any polynomial graph filter $H$. Consequently, filtering operations must be performed in the node domain instead of the frequency domain. Nevertheless, we can still gain insight into designing graph filters in the node domain by examining their effects in the frequency domain.
@@
Informal Definition of the Problem: For each degree $K$, find the optimal polynomial filter $H_{K}$ with the corresponding polynomial $p_{K}$ such that $H_{K}(f)$ converges to $\pi(f){\bf 1}$ as quickly as possible. To achieve this convergence, the polynomial $p_{K}$ should attenuate the eigenvalues (or frequencies) other than 0 of $L$. Therefore, the polynomial graph filter should function as a low-pass filter.
3.2 Revisiting Birkhoff Ergodic Theorem and Problem Formulation
Note that the iteration in the ergodic theorem can be written as a polynomial of $P$ in the following form: $\frac{(\mathbb{Id}+P+\ldots+P^{t-1})}{t}\triangleq p_{t}(P),$ where $p_{t}(z)\triangleq\sum_{k=0}^{t-1}\frac{1}{t}z^{k}=\frac{1}{t}\frac{1-z^{t}}{1-z},\,\,\,z\in[-1,1).$ If we define the following polynomial $q_{t}(z)\triangleq p_{t}(1-z)$, then we can write the same expression in terms of $L$ as follows:
$p_{t}(P)$ $=q_{t}(L)=\frac{1}{t}\frac{1-(1-L)^{t}}{1-(1-L)}=\frac{1}{t}\frac{1-\sum_{k=0}^{t}{t\choose k}(-L)^{k}}{L}$
$=\sum_{k=1}^{t}\frac{1}{t}{t\choose k}(-1)^{k-1}L^{k-1}\triangleq\sum_{k=0}^{t-1}a_{e}(k)\,L^{k}.$
If we define the polynomial graph filter $H_{t}\triangleq q_{t}(L)$, then we observe that we are indeed applying the polynomial graph filter $H_{t}$ to the graph signal $f$ in each iteration of the ergodic theorem. As $t$ approaches infinity, this graph filter attenuates the frequencies different from zero. Specifically,
@@ 3.2 Revisiting Birkhoff Ergodic Theorem and Problem Formulation
for $z\in(0,2]$ and $q_{t}(0)=1$. This can be proved very easily since $p_{t}(z)\rightarrow 0$ as $t\rightarrow\infty$ for $z\in[-1,1)$. Consequently, $H_{t}(f)\rightarrow\pi(f){\bf 1}$. Hence, in each iteration of the ergodic theorem, we are applying a low-pass filter to the graph signal $f$. However, this low-pass filter is not the optimal one. The goal of this paper is to find the optimal low-pass filter in each iteration. This problem is very similar to the problem of designing optimal low-pass filters in classical signal processing [8]. Therefore, as in classical signal processing, we use results from approximation theory.
Given the non-zero eigenvalues $\{\lambda\}_{x\in{\mathsf{X}}\setminus\{x_{*}\}}\subset(0,2]$ of $L$, it is not the case that $q_{t}(\lambda_{x})=0$ for all $x\in{\mathsf{X}}\setminus\{x_{*}\}$ for sufficiently large $t$ values, where $q_{t}$ is the polynomial used in the ergodic theorem. However, if we have complete knowledge of the eigenvalues $\{\lambda\}_{x\in{\mathsf{X}}}$ of $L$, then we can use a Lagrange polynomial $q$ of degree $|{\mathsf{X}}|-1$ [22, Definition 6.1] to achieve:
@@
That is, we can filter out frequencies that are different from zero. However, in general, we do not have access to the full set of eigenvalues of $L$ and so constructing $q_{t}^{L}$ is unfeasible.
In general, the most that one can do is to estimate some important eigenvalues different than $\lambda_{x_{*}}$. For instance, we may obtain some lower bound to the second smallest eigenvalue of $L$ [6] [4, Section 9.2]. Indeed, if $\lambda_{x_{**}}$ denotes the second smallest eigenvalue of $L$, then $\lambda_{x_{**}}$ is equal to the spectral gap of $P$. In this case, if we have some lower bound on this second smallest eigenvalue: $\lambda_{x_{**}}\geq\lambda_{{\mathsf{low}}},$ then the best approach would be to design an ideal low-pass filter:
$q^{{\mathsf{ideal}}}(\lambda_{x_{*}})=1\,\,\text{and}\,\,q^{{\mathsf{ideal}}}(\lambda)=0\,\,\text{for}\,\,2\geq\lambda\geq\lambda_{{\mathsf{low}}}.$
However, achieving this is impractical because an ideal low-pass filter demands an infinite number of components, as highlighted in classical signal processing [22]. Therefore, rather than using an infinite-degree polynomial, we seek a polynomial of degree $t$ at each time step $t$ that meets the following criteria:
@@
Here, we also want to converge to zero for values in $\lambda_{{\mathsf{low}}},2]$ as quickly as possible while maintaining $q_{t}(\lambda_{x_{*}})=1$. In essence, we aim to closely approximate the ideal low-pass filter in the most efficient manner. To achieve this, we can view our problem as designing an optimal polynomial low-pass graph filter with a one-hop operator $L$ to eliminate frequencies above $\lambda_{{\mathsf{low}}}$, where $\lambda_{{\mathsf{low}}}$ is a fixed threshold. This problem shares strong parallels with the design of optimal FIR low-pass filters in classical signal processing [8]. Indeed, in our work, we rely on concepts and results from approximation theory that are quite similar to those used in optimal FIR low-pass filter design.
@@
$\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\triangleq\sup_{z\in[\lambda_{{\mathsf{low}}},2]}|p(z)|.$
Remark 3.8.
The optimization problem described above is precisely the same as the one underlying Chebyshev acceleration, also known as the Chebyshev semi-iterative method [21, 12, 17]. Mathematically, therefore, our optimal graph filter problem is equivalent to the filtering problem encountered in the Chebyshev semi-iterative method. This equivalence is not surprising, as such optimal filtering problems are well established in the numerical linear algebra and signal processing literature. Indeed, the solution is given by a classical result from approximation theory concerning Chebyshev polynomials. Consequently, from a purely mathematical standpoint, this problem has already been extensively treated across various communities, beginning with approximation theory and extending to signal processing. We therefore make no claim of novelty in this regard; we are simply applying well-known results from approximation theory. The same holds for the other filter designs considered in this paper (Bernstein and Legendre).
Our contribution lies instead in the novel perspective we bring to this problem through the lens of graph signal processing. This reinterpretation is significant because it opens the door to applying a broader range of well-established filtering techniques, such as IIR filters, to further accelerate convergence. Moreover, by framing the problem within signal processing, one can now draw upon a rich set of tools and methodologies from that field to address related challenges in this setting.
Let us now give a motivation for this optimization problem. Note that, originally, we want to minimize $\|p_{t}(L)f-\pi(f){\bf 1}\|_{\pi}$ for any unit norm graph signal $f$, where $p_{t}\in P_{t-1}[\lambda_{{\mathsf{low}}},2]$ with $p_{t}(0)=1$, for each time step $t\geq 2$ as any graph signal is viewed as an element of the inner product space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$ [^1] In this context, we can use any norm we prefer (for instance, sup-norm) because the interpretation at the end will be the same.. An upper bound to this term can be obtained as follows:
$\|p_{t}(L)f-\pi(f){\bf 1}\|_{\pi}=\left\|\sum_{x\in{\mathsf{X}}}{\hat{f}}(x)p_{t}(\lambda_{x})f_{x}-{\hat{f}}(x_{*})f_{x_{*}}\right\|_{\pi}=\left\|\sum_{x\neq x_{*}}{\hat{f}}(x)p_{t}(\lambda_{x})f_{x}\right\|_{\pi}$
$\leq\sum_{x\neq x_{*}}|{\hat{f}}(x)||p_{t}(\lambda_{x})|\left\|f_{x}\right\|_{\pi}=\sum_{x\neq x_{*}}|{\hat{f}}(x)||p_{t}(\lambda_{x})|\leq\left(\sum_{x\neq x_{*}}|{\hat{f}}(x)|\right)\sup_{x\neq x_{*}}|p_{t}(\lambda_{x})|$
In the last inequality above, the first term depends on the graph signal $f$, and so, we cannot control this. Thus, our objective should be to minimize the second term by choosing $p_{t}$ appropriately. However, because we do not have access to the full eigenvalue structure, we cannot directly minimize the second term. Given the lower bound $\lambda_{{\mathsf{low}}}$ for the second smallest eigenvalue of $L$, we can instead try to minimize $\|p_{t}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}$, which leads to the optimization problem $\bf(OPT_{t})$}.
4 Filter Design with Bernstein Polynomials
For each time step $t\geq 2$, we want to design a polynomial $p$ of degree at most t1t-1 so that